# Recurrence Criteria for Reducible Homogeneous Open Quantum Walks on the Line

### v2

@@ 4 A General Recurrence Criteria for Open Quantum Walks
Following the approach described in [6, Section 4], as well as the references therein, we can obtain for a coin $(L,R)$ a decomposition
$\mathfrak{h}=\mathfrak{R}\oplus\mathfrak{T},\qquad\mathfrak{R}=\sup\{\mbox{supp}(\omega)\;|\;\omega\mbox{ is an invariant state for }\mathcal{L}\}.$
@@
Further, every $\mathfrak{X}_{\alpha}$ either is a minimal enclosure or assumes a decomposition (not unique) as the sum of mutually orthogonal isomorphic minimal enclosures $\{\mathfrak{V}_{\alpha,\beta},\;\beta\in I_{\alpha}\}:$
$\mathfrak{X}_{\alpha}=\bigoplus_{\beta\in I_{\alpha}}\mathfrak{V}_{\alpha,\beta}.$
Given a minimal enclosure $\mathfrak{V}_{\beta},$ let $\tau_{\beta}$ be the unique invariant state with support on $\mathfrak{V}_{\beta}.$ Introducing the parameter
@@
we have by [6, Lemma 4.2] that $m_{\beta}$ is independent of the particular minimal enclosure $\mathfrak{V}_{\beta}$ in $\mathfrak{X}_{\alpha}.$ Therefore, we can combine Proposition 3.4, Theorem 3.7 and the discussion above to state the property
@@
We remark that the convergence in law brought by corollary 4.2 can be upgraded to a convergence a.s., as demonstrated in [9, Theorem 3.5.2].
Let us start a HOQW induced by a coin $(L,R).$ There are two possibilities concerning recurrence when $\text{supp}(\rho_{0})\not\perp\mathcal{X}:$
1. $(L,R)$ is recurrent for $\mathcal{Y}_{\alpha}$ for some $\alpha$ and supp$\left(\mathcal{L}^{n}(\rho_{0})\right)\not\perp\mathcal{Y}_{\alpha}$ for some $n\in\mathbb{N}$: in this case, the walk is $\rho_{0}$-recurrent. Indeed, concerning the quantum trajectories $(\rho_{n})_{n}$, we could refresh the walk on some instant $n$ at some vertex $\lvert X_{n}\rangle,$ which is $\rho_{n}$-recurrent, since supp$\left(\mathcal{L}^{n}(\rho_{0})\right)\not\perp\mathcal{Y}_{\alpha}$ for a sufficiently large $n$. If there was some density $\sigma$ with supp$(\sigma)\not\perp\mathcal{X}$ and $\sigma$-transient, there would be some enclosure not orthogonal to $\mathcal{X},$ which is a contradiction by Proposition 3.4. The conclusion is held by the homogeneity of the coin, since, in this case, the walk either is $\rho_{s}$-recurrent for every $s=0,1,2,\ldots$ or it is $\rho_{s}$-transient for every $s=0,1,2,\ldots$, analogously to the proof of Proposition 3.4.
2. $(L,R)$ is transient for all $\mathcal{Y}_{\alpha}$ satisfying supp$\left(\mathcal{L}^{n}(\rho_{0})\right)\not\perp\mathcal{Y}_{\alpha}$ for every $n\in\mathbb{N}$: the walk is $\sigma$-transient for all $\sigma\in\mathcal{D}(\mathcal{X})$. Indeed, in this case, starting the walk on vertex $X_{0}=0,$ $X_{n}/n$ converges (when $n\rightarrow\infty$) to
@@
as given by Equation (6), where the second equality follows from $m_{\alpha}\neq 0,$ since $(L,R)$ is $\mathcal{Y}_{\alpha}$ transient. Thus, $X_{n}$ converges either to $\infty$- ∞ or $\infty$,∞ , in probability, imposing that the walk is $\rho$-transient and thus $(L,R)$ is transient on $\mathcal{D}(\mathcal{X}).$
@@
$(1)\quad\mathrm{Tr}(L\tau_{j}L^{*})\neq\dfrac{1}{2}\;\forall j\in\{1,\ldots m\}\Rightarrow(L,R)\mbox{ is transient;}$;
$(2)\quad\mathrm{Tr}(L\tau_{j_{k}}L^{*})=\dfrac{1}{2}\;$ for a set of indices $j_{k}\in J\subset\{1,\ldots,m\}$ and $\mathrm{Tr}(L\tau_{j_{l}}L^{*})\neq\dfrac{1}{2}\;$ for the remaining indices $j_{l}\in J^{C},$ and $\sigma$ is a density such that supp$(\sigma)\not\perp\mathcal{X}.$
(2.1) supp$\left(\mathcal{L}^{n}(\sigma)\right)\not\perp\mathcal{Y}_{\alpha}$ for some $\alpha\in J$ and some $n\in\mathbb{N},$ then $(L,R)$ is
@@
(2.2) supp$\left(\mathcal{L}^{n}(\rho)\right)\perp\mathcal{Y}_{\alpha}$ for all $\alpha\in J$ and all $n\in\mathbb{N},$ then $(L,R)$ is
@@
Remark 4.4.
It is worth observing that $\mathcal{D}_{2}$ in Theorem 4.3 actually arises from an enclosure due to the orthogonality between supp$\left(\mathcal{L}^{n}(\rho)\right)$ and $\mathcal{Y}_{\alpha}$; otherwise, it would contradict Proposition 3.4.
5 Examples
@@
    $\mathrm{Tr}(L\rho_{\infty}L^{*})=\frac{n^{2}}{2}=\mathrm{Tr}(R\rho_{\infty}R^{*})\Rightarrow\;(L,B,R)\mbox{ is recurrent,}$,
@@
    $(1)\quad|x_{1}|=\frac{1}{\sqrt{3}}\mbox{ and }|x_{2}|=\frac{1}{2}$$\Rightarrow$ $(L,B,R)\mbox{ is recurrent,}$,
    $(2)\quad|x_{1}|\neq\frac{1}{\sqrt{3}}\mbox{ and }|x_{2}|\neq\frac{1}{2}$$\Rightarrow$ $(L,B,R)\mbox{ is transient,}$,
    $(3)\quad|x_{1}|=\frac{1}{\sqrt{3}}\mbox{ and }|x_{2}|\neq\frac{1}{2}$$\Rightarrow$ $(L,B,R)\mbox{ is transient with respect to }\sigma=\begin{bmatrix}0&0\\ 0&1\end{bmatrix}\mbox{ and it is}$
    recurrent with respect to all densities but $\mbox{recurrent with respect to all densities but }\sigma,$
    $(4)\quad|x_{1}|\neq\frac{1}{\sqrt{3}}\mbox{ and }|x_{2}|=\frac{1}{2}$$\Rightarrow$ $(L,B,R)\mbox{ is transient with respect to }\sigma=\begin{bmatrix}1&0\\ 0&0\end{bmatrix}\mbox{ and it is}$
    recurrent with respect to all densities but $\mbox{recurrent with respect to all densities but }\sigma.$
@@
We remark that the choices of the coins above were taken in such a way that $L,B^{\prime}$ and $R$ had a common invariant eigenvector, namely $[1,0]^{T},$ which was not an eigenvector of $B.$ This is due to the hard job of, given non-trivial $R$ and $L,$ to find $B^{\prime}$ in such a way that $(L,B^{\prime},R)$ is recurrent. Attempt were conduct using mathematic software with many $L$’s and $R$’s having simple entries, but even though the entries of $B$’s that made $(L,B,R)$ to be transient were very long to write in this paper.
@@
Let us consider the coin given in [6, Section 6.2]:
@@
Consider a HOQW induced by the coin $(L,R),$ where
@@
6 Conclusion
We present new general recurrence criteria for classes of HOQWs, whose dynamics are governed by Kraus operators that are induced by a finite-dimensional coin.
In the context of these HOQWs, we presented three distinct criteria for analyzing recurrence. Firstly, Theorem 3.7 introduces a criterion for ergodic Lazy coins, where the recurrence is determined by the operators $L$ and $R$ and by the only invariant state $\rho_{\infty}$ associated with the corresponding auxiliary map.
Next, Theorem 3.9 focuses on Lazy HOQWs of dimension 2, offering a complete and easy-to-apply criterion. The simplicity of this criterion comes from the fact that it depends on the knowledge of the eigenvectors of the operators $L$, $B$, and $R$, allowing a practical analysis of the recurrence of these processes that are not necessarily ergodic.
Finally, Theorem 4.3 extends the analysis to non-lazy coins, using a decomposition of the vector space where the operators $L$, $B$, and $R$ act. By checking recurrence in vector subspaces that contain unique invariant states, this general criterion is extended to any initial density, significantly expanding the scope of recurrence analysis for HOQWs in more complex configurations, since it is valid for any finite-dimensional coin.

### v3

@@ 4 A General Recurrence Criteria for Open Quantum Walks
Following the approach described in [10, Section 4], as well as the references therein, we can obtain for a coin $(L,R)$ a decomposition
[…]
@@
Further, every $\mathfrak{X}_{\alpha}$ either is a minimal enclosure or assumes a decomposition (not unique) as the sum of mutually orthogonal isomorphic minimal enclosures $\{\mathcal{Y}_{\alpha,\beta},\;\beta\in I_{\alpha}\}:$
$\mathfrak{X}_{\alpha}=\bigoplus_{\beta\in I_{\alpha}}\mathcal{Y}_{\alpha,\beta}.$
Given a minimal enclosure $\mathcal{Y}_{\beta},$ let $\tau_{\beta}$ be the unique invariant state with support on $\mathcal{Y}_{\beta}.$ Introducing the parameter
@@
we have by [10, Lemma 4.2] that $m_{\beta}$ is independent of the particular minimal enclosure $\mathcal{Y}_{\beta}$ in $\mathfrak{X}_{\alpha}.$ Therefore, we can combine Proposition 3.5, Theorem 3.8 and the discussion above to state the property
@@
We remark that the convergence in law brought by corollary 4.2 can be upgraded to a convergence a.s., as demonstrated in [13, Theorem 3.5.2].
Let us start an HOQW induced by a coin $(L,R).$ There are two possibilities concerning recurrence when $\text{supp}(\rho_{0})\not\perp\mathcal{X}:$
[…]
@@
as given by Equation (6), where the second equality follows from $m_{\alpha}\neq 0,$ since $(L,R)$ is $\mathcal{Y}_{\alpha}$ transient. Thus, $X_{n}$ converges either to -$\infty$ or $\infty,$ in probability, imposing that the walk is $\rho$-transient and thus $(L,R)$ is transient on $\mathcal{D}(\mathcal{X}).$
@@
$(1)\quad\mathrm{Tr}(L\tau_{j}L^{*})\neq\dfrac{1}{2}\;\forall j\in\{1,\ldots m\}\Rightarrow(L,R)\mbox{ is transient;}$
[…]
@@
[…]
@@
Remark $4.$
[…]
5 Examples - 1D
@@
    $\mathrm{Tr}(L\rho_{\infty}L^{*})=\frac{n^{2}}{2}=\mathrm{Tr}(R\rho_{\infty}R^{*})\Rightarrow\;(L,B,R)\mbox{ is recurrent,}$
@@
    $(1)\quad|x_{1}|=\frac{1}{\sqrt{3}}\mbox{ and }|x_{2}|=\frac{1}{2}$ $\Rightarrow$ $(L,B,R)\mbox{ is recurrent,}$
    $(2)\quad|x_{1}|\neq\frac{1}{\sqrt{3}}\mbox{ and }|x_{2}|\neq\frac{1}{2}$ $\Rightarrow$ $(L,B,R)\mbox{ is transient,}$
[…]
    $\mbox{recurrent with respect to all densities but }\sigma,$
[…]
    $\mbox{recurrent with respect to all densities but }\sigma.$
@@
We remark that the choices of the coins above were taken in such a way that $L,B^{\prime}$ and $R$ had a common invariant eigenvector, namely $1,0]^{T},$ which was not an eigenvector of $B.$ This is due to the hard job of, given non-trivial $R$ and $L,$ to find $B^{\prime}$ in such a way that $(L,B^{\prime},R)$ is recurrent. Attempt were conduct using mathematic software with many $L$’s and $R$’s having simple entries, but even though the entries of $B$’s that made $(L,B,R)$ to be transient were very long to write in this paper.
@@
Let us consider the coin given in [10, Section 6.2]:
@@
Consider an HOQW induced by the coin $(L,R),$ where
@@
6 2D Open Quantum Walks
We start this section with the continuous-time OQW.
6.1 Continuous-Time Open Quantum Walks
In order to define the CTOQWs, we recall that an operator semigroup $\mathcal{T}$ on a Hilbert space $\mathcal{K}$ is a family of bounded linear operators $(T_{t})$ acting on $\mathcal{K}$, $t\geq 0,$ such that
$T_{t}T_{s}=T_{t+s},\;\forall s,t\in\mathbb{R}^{+}\;\mbox{ and }\;T_{0}=I_{\mathcal{K}}.$
If $t\mapsto T_{t}$ is continuous for the operator norm of $\mathcal{K}$, then $\mathcal{T}$ is said to be uniformly continuous, which is equivalent to the following assertion [7]: There exists a bounded linear operator $L$ on $\mathcal{K}$ such that $T_{t}=e^{tL},\;\forall t\geq 0.$ In this case,
$L=\lim_{t\rightarrow\infty}\frac{1}{t}(T_{t}-I_{\mathcal{K}})$
and the operator $L$ is called the generator of $\mathcal{T}.$